I need help solving the following systems. Please show the work so that I can understand for future reference. Thanks in advance for any help.
x+3/2y=7/2
3x=3/4y

Answer by J2R2R(94)

x + 3/2y = 7/2
3x = 3/4y

From the second equation we have
x = y/4
So substituting this into the first equation we have
y/4 + 3y/2 = 7/2
x 4 y + 6y = 14
7y = 14
y = 2

Giving x = y/4 = 2/4 = 1/2

Check
1/2 + 3 = 7/2
3/2 = 6/4